OpenStudy (anonymous):

Choose the correct simplification of the expression a^12b^8/a^2b^4.

OpenStudy (anonymous):

@amistre64 @amistre64

OpenStudy (anonymous):

@blondie16

OpenStudy (anonymous):

these are polynomials ..

OpenStudy (anonymous):

well you get a^12 divided by a^2 = a^10 and b^8 divided by b^4=b^4 so your answer is a^10*b^4

OpenStudy (anonymous):

a^10b^4 right?